# Stabilizability and lower spectral radius for linear switched systems with singular matrices

### v2

@@ 3.3 A special case
It is helpful to consider a special case of System 1 by letting $a=2$ and $b=c=d=0$, i.e the following example
@@
Figure 1: Stabilizability radius for $\alpha\in(0,1)$
One vital property of Example 3.8 is that this system is pointwise stabilizable for most $\alpha$. Figure 1 shows the $\tilde{\rho}$ in terms of $\alpha$ for Example 3.8. Each point in figure with $\alpha$ as a sample from $U(0,1)$ and $\tilde{\rho}$ from Corollary 3.9. Notice that rational $\alpha$ with an even denominator gives us $\tilde{\rho}=0$ from Corollary 3.9, so we randomly smaple $\alpha$ in this way to avoid having too many zero values of $\tilde{\rho}$.
@@
For any $\alpha\in(0,1)$, we have $\tilde{\rho}\leq 1$ and the equality only hold when $\alpha=1/3\text{ or }2/3$.
@@
The Diophantine approximation of Example 3.8 is given by $R(\alpha)=\|l\alpha-1/2\|_{N}$. For irrational $\alpha$, $R(\alpha)$ could be arbitrarily small so there exists an $l_{0}$ such that $\|l_{0}\alpha-1/2\|_{N}<1/6$.
@@
Then we know for any $\alpha\in(0,1)\setminus\bigcup_{c\in[0,1)}L_{\beta}(c)$, Example 3.8 has the finiteness property w.r.t. $\tilde{\rho}$, and each $L_{\beta}(c)$ has zero Hausdorff dimension. However the set $\bigcup_{c\in[0,1)}L_{\beta}(c)$ is not necessarily a zero Hausdorff dimension set.
4 Some examples
From section $3$, we know that when $\alpha=1/3$, Example 3.8 is unstabilizable. However, we can obtain a stabilizable system from Example 3.8 by increasing the smallest singular value.
@@ 4 Some examples
Figure 2 shows the range of $\theta$ such that each of the above norm is less than $1$. One can obtain the precise range of $\theta$ by solving the inequalities that the norm is less than $1$, which is not shown here.
@@
We know that the vector is projected to a lower dimensional subspace after applying a singular matrix. One general way of constructing a optimal sequence is to focus on the movement of the image of singular matrices. For example, after applying $M_{2}$ once, Example 3.8 only describes movement of the one dimensional image in $\mathbb{R}^{2}$. For some singular matrices, we can minimize the dimension of the image first.
@@
In Example 4.3, the image of $M^{2}_{1}$ is a one dimensional subspace and the $M^{2}_{1}$ have the same largest singular value as $M_{1}$. To find the optimal sequence, we can first apply $M_{1}$ twice to reduce the dimension of the image, then apply $M_{2}$ to rotate the vector in the image subspace. Sometimes, matrix combinations can be used to reduce the dimension of the image.
@@
We know the image of the matrix combination $M_{1}M_{2}M_{1}$ in Example 4.4 is a one dimensional subspace. To stabilize this system, we can apply $M_{1}M_{2}M_{1}$ to any initial vectors first. However, maximizing the dimension of kernel by some matrix sequence is not always the best strategy to find the optimal sequence.
@@ 5 Conclusion
In this paper, two different approaches to understand the stabilizability radius of linear discrete time switching systems are presented. First, we showed that in general, the stabilizability radius is lower bounded by the joint spectral subradius divided by the number of matrices. This result leads to more relationships between the stabilizability radius and joint spectral subradius. Then we focused on the two dimensional switched system with a singular matrix and a matrix with complex eigenvalues, in which we presented a method to compute the exact stabilizability radius. The parameter set is a zero Hausdorff dimension set when the stabilizability radius equals a constant, such as 0, are presented using this method.
Although most works of linear discrete time switching systems in the literature focus on general results, we presented some examples to show systems with singular matrices are also interesting. The singular matrices function as projections in systems, which could sometimes simplify the optimal sequence of matrices. However, it could also lead to more complex behaviors. For instance, in Example 3.8, the set of $\alpha$ such that the stabilizability radius equals 0 has zero Hausdorff dimension and dense in $(0,1)$. In Example 4.5, we need to consider more matrix combinations to find the optimal sequence.
There are still many open questions about the stabilizability radius of linear discrete time switching systems. Such as finding a more precise lower bound of the stabilizability radius, or give more situations when the stabilizability radius equals the joint spectral subradius. In particular, for switched systems with singular matrices, could we find a general method to decide when would stabilizability radius equals zero? Can we say more about the finiteness property w.r.t. the stabilizability radius of System 1 or other systems?
@@ Acknowledgments
The authors are grateful to Alan Haynes for pointing them to Ref. [1]. For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.

### v3

@@ 6 More low dimensional examples
It is helpful to consider a special case of System 4.1 by letting $a=2$ and $b=c=d=0$, i.e the following example.
@@
Figure 1: Stabilizability radius for $\alpha\in(0,1)$ of Example 6.1
One property of Example 6.1 is that this system is pointwise stabilizable for most $\alpha$. Figure 1 shows the $\tilde{\rho}$ in terms of $\alpha$ for Example 6.1. To generate this figure, choose $\alpha$ randomly sampled from $U(0,1)$ and calculate $\tilde{\rho}$ from Corollary 6.2. If only rational $\alpha$ are chosen, all $\tilde{\rho}$ values would be zero.
@@
For any $\alpha\in(0,1)$, we have $\tilde{\rho}\leq 1$ and the equality only holds when $\alpha=1/3\text{ or }2/3$.
@@
The Diophantine approximation of Example 6.1 is given by $R(\alpha)=\|l\alpha-1/2\|_{N}$. For irrational $\alpha$, $R(\alpha)$ could be arbitrarily small, so there exists an $l_{0}$ such that $\|l_{0}\alpha-1/2\|_{N}<1/6$.
@@
In this example, subset $S$ in Corollary 4.11 has Lebesgue measure close to $1$ if $c_{0}$ is close to $1$.
We know that when $\alpha=1/3$, Example 6.1 is unstabilizable. However, we can obtain a stabilizable system from Example 6.1 by increasing the smallest singular value.
@@
Figure 2 shows the range of $\theta$ such that each of the above norms is less than $1$. One can obtain the precise range of $\theta$ by solving the inequalities that the norm is less than $1$, which is not shown here.
@@
We see that the vector is projected to a lower dimensional subspace after applying a singular matrix. One general way of constructing a optimal sequence is to focus on the movement of the image of singular matrices. For example, after applying $M_{2}$ once, Example 6.1 only describes movement of the one dimensional image in $\mathbb{R}^{2}$. For some singular matrices, we can minimize the dimension of the image first.
@@
In Example 6.6, the image of $M^{2}_{1}$ is a one dimensional subspace and $M^{2}_{1}$ has the same largest singular value as $M_{1}$. To find the optimal sequence, we can first apply $M_{1}$ twice to reduce the dimension of the image, then apply $M_{2}$ to rotate the vector in the image subspace. Sometimes, matrix combinations can be used to reduce the dimension of the image.
@@
The image of the matrix combination $M_{1}M_{2}M_{1}$ in Example 6.7 is a one dimensional subspace. To stabilize this system, we can apply $M_{1}M_{2}M_{1}$ to any initial vector first. However, maximizing the dimension of kernel by some matrix sequence is not always the best strategy to find the optimal sequence.
@@ 7 Conclusion
In this paper, two different approaches to understand the stabilizability radius of linear discrete time switched systems with singular matrices are presented. First, we showed that in general, the stabilizability radius is lower bounded by the joint spectral subradius divided by the number of matrices. We also show the stabilizability radius equals joint spectral subradius for switched system with rank one matrices.
Then we focused on the two dimensional switched system with a singular matrix and an arbitrary matrix, in which we presented a method to compute the exact stabilizability radius. For two dimensional switched system with a singular matrix and a matrix with complex eigenvalues, the parameter set is a zero Hausdorff dimension set when the stabilizability radius equals a constant, such as 0, are presented using this method.
Although most works of linear discrete time switching systems in the literature focus on general results, we presented some examples to show systems with singular matrices are also interesting. The singular matrices function as projections in systems, which could sometimes simplify the optimal sequence of matrices. However, it could also lead to more complex behaviors. For instance, in Example 6.1, the set of $\alpha$ such that the stabilizability radius equals 0 has zero Hausdorff dimension and is dense in $(0,1)$. In Example 6.8, we need to consider more matrix combinations to find the optimal sequence.
There are still many open questions about the stabilizability radius of linear discrete time switching systems, such as finding a more precise lower bound of the stabilizability radius, or giving more situations when the stabilizability radius equals the joint spectral subradius. In particular, for switched systems with singular matrices, could we find a general method to decide when the stabilizability radius equals zero? Can we say more about the finiteness property w.r.t. the stabilizability radius of System 4.1 or other systems?
@@ Acknowledgments
The authors are grateful to Alan Haynes for pointing them to Ref. [1]. For the purpose of open access, the authors has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.